# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v2

@@ IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the number $L$ and the length $N$ of the chains, and contribute to reveal the inter-chain joint effect, to find optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12. It can be seen that for the three configurations SSH ST, TC-SSH ST, and TC-SSH TT the number $L$ of the chains has obvious effect in $L\leq 6$ [Fig. 11(e, f, h)]. Although there is large input exciton current, the occupation number $p_{k,N}$ of the last column is so few that the outgoing exciton current is little for SSH TT and the excitons are aggregated on the inside of the multichains with $L$ for SSH ST [Fig. 11(c1, e1, g1, d1)]. Therefore SSH ST is perfect candidate applied to quantum battery in the six configurations. What’s more, for SSH TT the most of excitons is occupied by the upper edge sites, which is similar to skin effect. In fact, excitons is restricted on the first chain with $L$ because of the low efficient transport of excitons among chains. On the contrary, because of the expand of excitons with $L$ the inner occupation number $p_{\rm bulk-in}$ increase and the occupation number of the first chain decrease for SSH ST. In the presence of cavity the steady-state dynamics display abundant phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinct oscillation with $L$ for SSH TT and TC-SSH TT, in which $I_{\rm i,o}$ and $\eta$ are peak (trough) for odd (even) number of $L$ and $N=10$ [Fig. 11(c2, g2, d2)]. The difference between steady-state dynamics of odd and even number become small with $L$. In addition to, the exciton transport is strengthened by cavity for SSH ST.
@@
There is an obvious character that the exciton transport efficiency $\eta$ behaves oscillation with the length $N$ of the chains in Fig. 12. We firstly discuss the situation without cavity in Fig. 12(a-d). The oscillation is relative weak for TC ST and depends on the parity of $L$ and $N$ for TC TT in the absence of cavity. Thus the diagonal inter-chain coupling induces difference between exciton transports of odd and even $N$ in long TC TT multichains as shown in Fig. 12(b,d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for SSH ST and SSH TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing for later. However, in the mixed multichains of TC-SSH because of the presence of TC chains the dimerization is suppressed, the oscillation is enhanced for TC-SSH ST and TC-SSH TT. In the two configurations the oscillation is consistent for odd $L$ and is opposite for even $L$ each other about the parity of $N$, such as $\eta$ of $N=16$ is trough and that of $N=17$ is peak for TC-SSH TT when $L=2$, but it is inverse for TC-SSH ST. When the cavity is considered the oscillation is significantly strengthened for the configurations containing SSH chain, as shown in Fig. 12(e-h). For SSH ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for SSH ST and SSH TT is consistent. Interestingly, it behaves inverse dependence on the parity of $L$ between SSH TT and TC-SSH ST, TC-SSH TT, i.e. for the former the oscillation amplitude of odd $L$ is bigger than that of even $L$, and for the latter the oscillation amplitude of even $L$ is bigger than that of odd $L$. Furthermore, the dependence of the exciton transport efficiency on the parity of $N$ can be changed by $L$ and cavity for TC-SSH ST and TC-SSH TT. For example, comparing with $L=2$ and $L=3$ for TC-SSH ST $N=16$, the $\eta$ corresponds respectively to peak and trough [Fig. 12(a, b)], as well as comparing with Fig. 12(b) and (f) for TC-SSH ST $N=14$ it corresponds respectively to trough and peak. We also note by comparing with Fig. 12(b) and (f), as well as $L=3$ [Fig. 12(f)] and $L=4$ [Fig. 12(g)] for TC-SSH ST, that the period of oscillation with $N$ also is changed by cavity and $L$.
@@ V Summary
In summary, the steady-state dynamics of exciton in the hybrid system consist of a 2D multichains and a cavity are studied through the spectra and the Lindblad quantum master equation, in which six inter-TLS coupling configurations are chosen as TC ST, TC TT, SSH ST, SSH TT, TC-SSH ST and TC-SSH TT based on the TC and SSH model, as well as square and triangle type. The effects of the inter-chain coupling strength, the dimerization parameter, the number of chain, and the length of chain on the spectra and the steady-state dynamics are investigated. The results show that in the absence of cavity the exciton transport efficiency is decided by the distribution of population of exciton on the whole multichains. Besides, when the cavity is considered the occupation number of the cavity photons depend significantly on the cavity-dressed energy level crossings and anticrossings near the zero-energy modes, at which the coherent excitation and LZ transition between the multichains and the cavity occur. Therefore, the exciton transport efficiency and currents via the polariton channel are changed at the locations of the crossings and anticrossings. We can conclude that the exciton transport dynamics are controlled by the crossings and anticrossings. The spectra display that the locations of the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter, as well as the stead-state dynamics also are correspondingly. Particularly, the exciton transport efficiency and currents are symmetric about the inter-chain coupling strength $\zeta=0$ in the multi-chain system without cavity, but the symmetry can be broken by cavity. In the fully dimerized limit, the presence of cavity enhances exciton transport for SSH ST, TC-SSH ST and TC-SSH TT of the even length of the chains. Interestingly, when the cavity drive is simultaneously exerted, the exciton transport is not limited by the occupation number of site $11$ and is strengthened for SSH ST and TC-SSH ST of the odd length of the chains in the negative range of the dimerization parameter. Further, in the space of the TLS transition frequency detuning and the inter-chain coupling strength we discover that the exciton transport is suppressed for TC ST and TC TT and strengthened for the other configurations by cavity at the range corresponding to the crossings and anticrossings, in which the polaritons act as vital role and the effects are enhanced by the cavity drive. In the space of the TLS transition frequency detuning and the dimerization parameter, the steady-state dynamics depend on the coupling configuration, driven method, and cavity, especially in the zone of the fully dimerized limit. Finally, it can be found that the exciton transport is obviously related to the parity of the number and length of the chains with cavity, in which the exciton transport efficiency behaves oscillation with them. The discrepancy between of the efficiency and currents for odd and even is related to the configuration. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like topological phase transition of the system.

### v3

@@ IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the number $L$ and length $N$ of the chains, and contribute to reveal the inter-chain joint effect, to find optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12. It can be seen in Fig. 11(c1111, g1111, d1111) that in the absence of the cavity the number $L$ of chains has obvious effect on exciton transport of the three configurations SSH ST, TC-SSH ST, and TC-SSH TT for $L\leq 6$. For HE TT, although the input exciton current is large, the occupation number $p_{k,N}$ of the last column limits the outgoing exciton current, leading to that $L$ has not effect on exciton transport. In fact, because of the inefficient transport of excitons among chains for HE TT, the most of excitons locate at the upper edge sites [Fig. 11(b1)], which is similar to skin effect. On the contrary, because of the diffusion of excitons among multichains the inner occupation number $p_{\rm bulk-in}$ increase and the occupation number of the edge chains decrease with $L$ for HE ST as shown in Fig. 11(b1, e1, f1). Therefore, for HE ST the excitons are aggregated on the inside of the multichains with large $L$ [Fig. 11(e1)], thus it is perfect candidate applied to quantum battery in the six configurations. In the presence of the cavity the steady-state dynamics represent abundant phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinct periodic oscillation with $L$ for HE TT and HO-HE TT, in which the peaks (troughs) of $I_{\rm i,o}$ and $\eta$ correspond to the odd (even) numbers of $L$ for $N=10$ [Fig. 11(c2, g2, d2)]. The result is contrary to that of single chain. The difference between the steady-state dynamics of the odd and of the even number becomes small with $L$. In addition to, the exciton transport is strengthened by the cavity for HE ST and HE TT.
@@
There is an obvious character that the exciton transport efficiency $\eta$ behaves oscillation with the length $N$ of the chains in Fig. 12. We firstly discuss the situation without cavity in Fig. 12(a-d). The oscillation is weak for HO ST relative to the other configurations and depends on the parity of $L$ and $N$ for HO TT in the absence of cavity. Thus the diagonal inter-chain coupling induces difference between exciton transports of odd and of even $N$ in long HO TT multichains as shown in Fig. 12(b, d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for HE ST and HE TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing on the multichains for later. However, in the mixed long multichains of HO-HE because of the presence of HO chains the dimerization is suppressed, the oscillation is enhanced for HO-HE ST and HO-HE TT. For the two configurations the oscillations with $N$ are consistent for odd $L$ and are opposite for even $L$ each other, such as when $L=3$ and $N=16$ $\eta$ is trough for both HO-HE TT and HO-HE ST, while $L=4$ and $N=16$ $\eta$ is peak for HO-HE ST and is trough for HO-HE TT. When the cavity is considered the oscillation is significantly strengthened for the configurations containing HE chain, as shown in Fig. 12(e-h). For HE ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for HE ST and HE TT is consistent, which illustrates further that the oscillation is mainly caused by HE chain. Interestingly, the oscillation amplitude behaves inverse dependence on the parity of $L$ between HE TT and HO-HE ST, HO-HE TT, i.e. for the former it with odd $L$ is bigger than that with even $L$, and for the latter it with even $L$ is bigger than that with odd $L$. Furthermore, the dependence of the exciton transport efficiency for HO-HE ST and HO-HE TT on the parity of $N$ can be changed by $L$ and the cavity. For example, comparing $L=2$ and $L=3$ for HO-HE ST of $N=16$, $\eta$ corresponds respectively to peak and trough [Fig. 12(a, b)], as well as comparing Fig. 12(b) and 12(f) for HO-HE ST of $N=14$ and $L=2$ it corresponds respectively to trough and peak. We also note by comparing Fig. 12(b) and 12(f), Fig. 12(f) and 12(g) for HO-HE ST, that the period of the oscillation with $N$ also is varied by the cavity and $L$.
@@ V Summary
In summary, in order to reveal the mechanism of exciton transport in a 2D multichains system with a cavity, the spectra, the second order partial derivative of the Hopfield coefficients of cavity photons with respect to the inter-chain coupling strength and the dimerization parameter, the von Neumann entropy and the steady-state dynamics of exciton are obtained by numerically diagonalizing the Hamiltonian and solving the Lindblad quantum master equation, in which six inter-TLS coupling configurations are chosen as HO ST, HO TT, HE ST, HE TT, HO-HE ST and HO-HE TT based on the HO and HE chain, as well as the square and triangle type of inter-chain coupling. The effects of the inter-chain coupling strength, the dimerization parameter, the detuning of the driving frequency to the TLS, the number of chain, the length of chain, and the coupling configurations on them are investigated. The results show that in the absence of cavity the exciton transport efficiency is decided by the distribution of population of exciton on the whole multichains. Besides, when the cavity is considered the occupation number of the cavity photons depend significantly on the cavity-dressed energy level crossings and anticrossings of the nearly-zero-energy modes, at which the coherent excitation and LZ transition between the multichains and the cavity occur. Therefore, the exciton transport efficiency and currents via the polariton channel are changed at the locations of the crossings and anticrossings. We can conclude that the exciton transport dynamics are controlled by the crossings and anticrossings. Besides, the locations of the crossings and anticrossings is accurately determined by calculating von Neumann entropy. We also discover that the second order partial derivative of the Hopfield coefficients of cavity photons is a good physical quantity to describe the locations of the crossings and anticrossings, and the change of photonic occupation number. The spectra show that the locations of the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter. Particularly, the exciton transport efficiency and currents are symmetric about the inter-chain coupling strength $\zeta=0$ in the multi-chain system without cavity, but the symmetry can be broken by the cavity. In the space of the dimerization parameter the presence of the cavity enhances exciton transport for HE ST, HO-HE ST and HO-HE TT with even chains, but suppresses it for HO-HE ST and HO-HE TT with odd chains when $\delta>0.36$. Interestingly, when the cavity drive is simultaneously exerted, the exciton transport is not limited by the occupation number of site $11$ and is strengthened for HE ST and HO-HE ST with odd chains in the range of the negative dimerization parameter. Further, in the space of the TLS transition frequency detuning and the inter-chain coupling strength we discover that the exciton transport is suppressed for HO ST and HO TT and is strengthened for the other configurations by the cavity at the range corresponding to the crossings and anticrossings, in which the polaritons act as vital role and the effects are enhanced by the cavity drive. In the space of the TLS transition frequency detuning and the dimerization parameter, the steady-state dynamics depend on the coupling configurations, the driven methods, and the cavity, especially in the zone of the fully dimerized limit. Finally, it can be found that the exciton transport is obviously related to the parity of the number and length of the chains with cavity, in which the exciton transport efficiency behaves oscillation with them. The oscillation is related to the configurations. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like the topological phase transition of the system.